Feedback

Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Summary

Feedback and Sensitivity in Climate

Alain Lahellec

11th June 2009

Feedback Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Summary

Outline

1 Motivation The Basic Problem That We Studied Previous Work

2 Our Results/Contribution Main Results Basic Ideas for Proofs/Implementation

Feedback Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Summary

1 Motivation The Basic Problem That We Studied Previous Work

2 Our Results/Contribution Main Results Basic Ideas for Proofs/Implementation

Feedback

Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Make Titles Informative. Use Uppercase Letters.

Subtitles are optional.

Use itemize a lot.

Use very short sentences or short phrases.

Feedback Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Summary

Make Titles Informative.

You can create overlays… using the pause command:

\{ \begin{matrix}∂_{t}σ_{0}(t) = M^{♭}σ_{0}(t, 0)+|b> Υ(t) \\ g(t) =<c| σ_{0}(t, 0) , \end{matrix}  (1)

Δθ(t) =<c(t)| Φ(t, 0)´ η_{0}^{r}

+ <c| (t) ∫_{0}^{t} Φ(t, τ)|b> (τ)Δθ_{0}(τ)dτ (2)

using overlay specifications:

First item. Second item.

with the definition of the propagator as:

∂_{t}Φ(t, τ) = M(t)Φ(t, τ) ; Φ(τ, τ) = I ,  (3)

with the identity matrix I. using the general uncover command:

First item. Second item.

Feedback Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Summary

You can create overlays… using the pause command:

\{ \begin{matrix}∂_{t}σ_{0}(t) = M^{♭}σ_{0}(t, 0)+|b> Υ(t) \\ g(t) =<c| σ_{0}(t, 0) , \end{matrix}  (1)

Δθ(t) =<c(t)| Φ(t, 0)´ η_{0}^{r}

+ <c| (t) ∫_{0}^{t} Φ(t, τ)|b> (τ)Δθ_{0}(τ)dτ (2)

using overlay specifications:

First item. Second item.

with the definition of the propagator as:

∂_{t}Φ(t, τ) = M(t)Φ(t, τ) ; Φ(τ, τ) = I ,  (3)

with the identity matrix I. using the general uncover command:

First item. Second item.

Feedback Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Summary

You can create overlays… using the pause command:

\{ \begin{matrix}∂_{t}σ_{0}(t) = M^{♭}σ_{0}(t, 0)+|b> Υ(t) \\ g(t) =<c| σ_{0}(t, 0) , \end{matrix}  (1)

Δθ(t) =<c(t)| Φ(t, 0)´ η_{0}^{r}

+ <c| (t) ∫_{0}^{t} Φ(t, τ)|b> (τ)Δθ_{0}(τ)dτ (2)

using overlay specifications:

First item. Second item.

with the definition of the propagator as:

∂_{t}Φ(t, τ) = M(t)Φ(t, τ) ; Φ(τ, τ) = I ,  (3)

with the identity matrix I. using the general uncover command:

First item. Second item.

Feedback Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Summary

You can create overlays… using the pause command:

\{ \begin{matrix}∂_{t}σ_{0}(t) = M^{♭}σ_{0}(t, 0)+|b> Υ(t) \\ g(t) =<c| σ_{0}(t, 0) , \end{matrix}  (1)

Δθ(t) =<c(t)| Φ(t, 0)´ η_{0}^{r}

+ <c| (t) ∫_{0}^{t} Φ(t, τ)|b> (τ)Δθ_{0}(τ)dτ (2)

using overlay specifications:

First item. Second item.

with the definition of the propagator as:

∂_{t}Φ(t, τ) = M(t)Φ(t, τ) ; Φ(τ, τ) = I ,  (3)

with the identity matrix I. using the general uncover command:

First item. Second item.

Feedback Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Summary

You can create overlays… using the pause command:

\{ \begin{matrix}∂_{t}σ_{0}(t) = M^{♭}σ_{0}(t, 0)+|b> Υ(t) \\ g(t) =<c| σ_{0}(t, 0) , \end{matrix}  (1)

Δθ(t) =<c(t)| Φ(t, 0)´ η_{0}^{r}

+ <c| (t) ∫_{0}^{t} Φ(t, τ)|b> (τ)Δθ_{0}(τ)dτ (2)

using overlay specifications:

First item. Second item.

with the definition of the propagator as:

∂_{t}Φ(t, τ) = M(t)Φ(t, τ) ; Φ(τ, τ) = I ,  (3)

with the identity matrix I. using the general uncover command:

First item. Second item.

Feedback Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Summary

You can create overlays… using the pause command:

\{ \begin{matrix}∂_{t}σ_{0}(t) = M^{♭}σ_{0}(t, 0)+|b> Υ(t) \\ g(t) =<c| σ_{0}(t, 0) , \end{matrix}  (1)

Δθ(t) =<c(t)| Φ(t, 0)´ η_{0}^{r}

+ <c| (t) ∫_{0}^{t} Φ(t, τ)|b> (τ)Δθ_{0}(τ)dτ (2)

using overlay specifications:

First item. Second item.

with the definition of the propagator as:

∂_{t}Φ(t, τ) = M(t)Φ(t, τ) ; Φ(τ, τ) = I ,  (3)

with the identity matrix I. using the general uncover command:

First item. Second item.

Feedback Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Summary

1 Motivation The Basic Problem That We Studied Previous Work

2 Our Results/Contribution Main Results Basic Ideas for Proofs/Implementation

Feedback

Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Summary

Figure: ClimSI WV feedback model responses.

Feedback

Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Feedback Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Summary

1 Motivation The Basic Problem That We Studied Previous Work

2 Our Results/Contribution Main Results Basic Ideas for Proofs/Implementation

Feedback

Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Feedback

Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Feedback

Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Feedback Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Summary

1 Motivation The Basic Problem That We Studied Previous Work

2 Our Results/Contribution Main Results Basic Ideas for Proofs/Implementation

Feedback

Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Feedback

Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Feedback

Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Feedback Author, Another

Motivation

The Basic Problem That We Studied Previous Work

Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Summary

Summary

The first main message of your talk in one or two lines. The second main message of your talk in one or two lines. Perhaps a third message, but not more than that.

Outlook

Something you haven’t solved. Something else you haven’t solved.

Feedback Author, Another

Appendix

For Further Reading

For Further Reading I

A. Author. Handbook of Everything. Some Press, 1990.

S. Someone. On this and that. Journal of This and That, 2(1):50–100, 2000.